OpenStudy (anonymous):

Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. What is the rate of the boat in still water? 2 mph 20 mph 22 mph

OpenStudy (gunsodo):

Let x represent the current's velocity... And let v represent the boat's velocity.... We'll get 6(v-x)=120 ------ 1 and 5(v+x)=120 ------------- 2 Simplify those equations v-x=20 -----------1 v+x=24 ---------- 2 Take 1 + 2 (v-x)+(v+x)=20+24 2v=44 v=22 The answer of this problem is 22 mph, or the third choice... Hope this help your question ^^

OpenStudy (anonymous):

yes thank you! @gunsodo